# The Open Proof Corpus: A Large-Scale Study of LLM-Generated Mathematical Proofs

### v1

@@ Abstract
In recent months, large language models (LLMs) have made significant progress in mathematical proof generation, but further advancement is hindered by the lack of a large-scale, high-quality dataset of human-evaluated proofs. While expensive to create, such a dataset is essential for driving improvements in training and enabling a rigorous analysis of proof generation capabilities. In this work, we present the Open Proof Corpus (OPC), a dataset comprising over 5,000 human-evaluated proofs produced by state-of-the-art LLMs. The OPC was specifically designed for broad applicability and downstream usage in proof generation research and is the first to include a substantial number of correct, LLM-generated solutions to problems from prestigious mathematics competitions such as the USAMO and IMO. Using the OPC, we explore critical questions in automated proof generation: (1) the performance gap between natural language and formal proof generation, (2) the discrepancy between final-answer accuracy and full-proof validity, and (3) the impact of best-of-n selection on proof quality. Finally, to showcase the utility of the OPC, we finetune an 8B-parameter model on the dataset, obtaining a model that performs on par with the best model, Gemini-2.5-Pro, on the task of evaluating proof correctness.
Machine Learning, ICML
Authors:
Dataset Contributors:
@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable progress in mathematical reasoning, attaining top-competitor performance on various final-answer benchmarks such as AIME and HMMT (Balunović et al., 2025). However, growing evidence suggests that these benchmarks fail to capture the full breadth of mathematical reasoning capabilities, as they do not require models to produce complete proofs or detailed intermediate steps (Mahdavi et al., 2025; Guo et al., 2025b). Such step-by-step reasoning is critical for applications in theorem proving, mathematical research, and education.
@@
To address this gap, several evaluation efforts have emerged, revealing that LLMs significantly underperform on proof generation compared to existing final-answer benchmarks (Petrov et al., 2025; Mahdavi et al., 2025; Guo et al., 2025b). Despite their value, these benchmarks are severely limited in their use for broader analysis, training, and future development purposes. Specifically, they are small in size (Petrov et al., 2025; Guo et al., 2025b), rely on outputs from outdated models (Frieder et al., 2023), contain few correct proofs (Mahdavi et al., 2025), and are not fully open-sourced (Mahdavi et al., 2025; Guo et al., 2025b).
@@
Furthermore, key questions about proof generation capabilities remain unanswered. First, while it is widely claimed that final-answer benchmarks are insufficient for evaluating proof generation capabilities, this claim is not yet substantiated by evaluating LLM-generated proofs on such benchmarks. Second, with recent advances in formal proof generation using systems like Lean (Ren et al., 2025; Wang et al., 2025), the performance gap between natural language and formal proof generation remains unclear. Third, the potential of best-of-n sampling strategies to improve proof quality has not been explored.
Sample Problem: EGMO 2012, Problem 3
Level: High School
Split: Generic
Question: Solve over $\mathbb{R}$ the functional equation $f(yf(x+y)+f(x))=4x+2yf(x+y)$ Refer to caption Proof: To solve the functional equation …we conclude that this is the **only** solution. Refer to caption Grade (Incorrect): The model substitutes $x$ to be 1. All further ideas are based on the wrong assumption. Refer to caption
@@
Figure 1: Overview of the OPC and its conclusions. On the left, we show a typical sample from the OPC, including a question taken from a high-quality mathematical competition, a proof generated by an LLM, and a human judgment of the proof’s correctness. On the right, we summarize the main conclusions drawn from the OPC.
@@
To address these challenges, we introduce the Open Proof Corpus (OPC): a large-scale, human-validated dataset comprising over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems. As shown in Fig.˜1(a), each OPC sample includes (1) a problem from a high-quality mathematical competition such as the International Mathematical Olympiad (IMO), (2) a proof generated by a state-of-the-art LLM, and (3) a binary human evaluation of the proof’s correctness with feedback. The OPC was specifically designed for downstream usage in proof generation research, enabling both training and evaluation of LLMs on proof generation tasks. To address the open questions outlined earlier, the OPC includes problems from specific sources: problems from the PutnamBench (Tsoukalas et al., 2024) enable comparisons between formal and informal reasoning, while problems from MathArena (Balunović et al., 2025) support evaluation of proof correctness for problems with final answers.
@@
Despite the difficulty of problems in the OPC, state-of-the-art LLMs demonstrate surprisingly strong performance. For instance, o4-mini correctly solves almost $20\%$ of the problems in the IMO Shortlist, and the OPC generally consists of $43\%$ correct proofs. Furthermore, as shown in the top left of Fig.˜1(b), LLMs exhibit strong capabilities in evaluating proofs: Gemini-2.5-Pro achieves $88.1\%$ accuracy in judging proof correctness—nearly matching human performance at $90.4\%$. To showcase the utility of the OPC, we fine-tune DeepSeek-R1-Qwen3-8B (Guo et al., 2025a) using GRPO (Shao et al., 2024) on the OPC, resulting in an open model that achieves $88.1\%$ judgment accuracy, on-par with Gemini-2.5-Pro and outperforming all other models.
@@
The OPC allows us to empirically resolve the open questions outlined above, with all conclusions shown in Fig.˜1(b). First, as shown on the top right, we find that natural language proof generation significantly outperforms formal proof generation: on PutnamBench, Gemini-2.5-Pro solves 11 times more problems than the best formal model, Deepseek-Prover-V2 (Ren et al., 2025). Second, as shown on the bottom left, we observe a substantial gap between final-answer accuracy and proof correctness. While Gemini-2.5-Pro loses only $8\%$ of its final-answer accuracy when proof correctness is required, o3 suffers a drop of almost $30\%$. Third, as shown on the bottom right, we find that best-of-n strategies yield large gains in performance. Interestingly, while standard best-of-n selection methods moderately improve accuracy from $26\%$ to $36\%$, a ranking-based approach inspired by (Liu et al., 2025) achieves the highest performance of $47\%$.
@@
  • A scalable pipeline for generating and evaluating natural language mathematical proofs (Section˜3).
  • The Open Proof Corpus, a large, human-validated dataset of over 5,000 LLM-generated proofs (Section˜4).
  • New benchmarking results on the OPC, resulting in novel insights into current LLM capabilities (Section˜5).
  • An open-source, 8B-parameter model fine-tuned on the OPC that achieves $88.1\%$ judgment accuracy, tied with the best model on this task (Section˜5).
@@ 2 Related Work
We briefly recapitulate the relevant literature on mathematical reasoning benchmarks and datasets.
@@
Final-answer benchmarks evaluate models by comparing a final answer from the model’s output with a ground-truth answer, typically using rule-based parsers. With the advent of reasoning LLMs (Jaech et al., 2024; Guo et al., 2025a), older benchmarks have become saturated (Cobbe et al., 2021; Lightman et al., 2024), and even more recent ones are nearing saturation (Gao et al., 2025; He et al., 2024; Balunović et al., 2025; Gulati et al., 2024). Only private benchmarks such as FrontierMath (Glazer et al., 2024) and HLE (Phan et al., 2025) remain sufficiently challenging for the latest models. However, these benchmarks do not require the generation of full proofs or detailed reasoning steps. Moreover, their private nature hinders reproducibility and broader community engagement, both key factors in driving progress.
@@
Another growing line of work involves training LLMs to generate formal proofs in languages such as Lean (de Moura and Ullrich, 2021) or Isabelle (Nipkow et al., 2002), which can then be automatically verified by these systems (Zheng et al., 2022; Tsoukalas et al., 2024; Yu et al., 2025). While this paradigm enables rigorous correctness checking, it typically requires models to be specifically finetuned for formal proof generation (Ren et al., 2025; Wang et al., 2025; Lin et al., 2025). In contrast, state-of-the-art general-purpose models like o3 and Gemini-2.5-Pro struggle with formal proof generation. As a result, formal reasoning currently does not allow to make full use of the natural language capabilities of these general-purpose models, and, as we show in Section˜5, there remains a significant performance gap between formal and natural language proof generation, with the latter being substantially more effective.
@@
Two concurrent works have recently expanded this line of research. Sheng et al. (2025) focus specifically on inequality proofs and primarily emphasize the development of an LLM as a judge framework to mitigate the high cost of human evaluation. Guo et al. (2025b) highlight a notable gap between final-answer accuracy and the ability to generate correct proofs, a finding we confirm in Section˜5. However, their analysis stops at this observation and does not evaluate performance on an established final-answer benchmark. Further, both studies are limited in scale compared to the OPC and have not open-sourced their human-annotated proofs.
@@
Several large-scale datasets have been developed to train LLMs on mathematical reasoning. One of the earliest, Li et al. (2024), compiled a large dataset of internet-sourced problems, including both final-answer questions and natural language proofs. However, it lacks LLM-generated proofs, human judgments, and examples of incorrect proofs. Other datasets have focused exclusively on final-answer tasks (Albalak et al., 2025; He et al., 2025; Moshkov et al., 2025), offering limited support for training or evaluating proof generation. Finally, Zhang et al. (2025) introduced a dataset of both valid and invalid problem statements, each accompanied by LLM-generated proofs. While this ensures that incorrect proofs exist for invalid statements, the dataset does not include human evaluation of proofs or any other information on the validity of the proofs.
LLM as a judge
The widespread success of LLMs has led to the development of the "LLM as a judge" paradigm, which enables scalable, consistent, and cost-effective evaluation of complex outputs (Zheng et al., 2023; Gu et al., 2024). This approach serves as a promising alternative to traditional expert-based evaluation and opens up new possibilities for self-assessment during generation through methods such as best-of-n sampling, feedback-guided decoding, and self-training (Pan et al., 2023; Madaan et al., 2023; Weng et al., 2023; Jiang et al., 2025). However, concerns remain about the robustness of this framework in high-stakes domains like mathematical reasoning, where issues such as overconfidence and self-degradation can undermine reliability (Huang et al., 2024; Pan et al., 2023).
@@ 3 Methodology
Accurately evaluating LLM-generated proofs is a challenging task. Models often make difficult-to-detect errors, and they rarely acknowledge when they cannot solve a problem (see Section˜5). In this section, we outline the methodology used to create the OPC, with particular emphasis on the complexities of evaluating LLM-generated proofs and our efforts to maximize dataset size. Since human judges cannot reasonably spend hours studying each problem, we developed a pipeline to support efficient grading. This pipeline consists of three main components: problem and judge preparation (Section˜3.1), the grading procedure (Section˜3.2), and monitoring and validation (Section˜3.3).
@@ 3.1 Problem and Judge Preparation
Judges were selected from among former IMO participants or individuals who reached the final stages of IMO selection in their respective countries. We contacted each judge personally to ensure they were qualified, motivated, and easily accessible for ongoing communication. A total of $13$ judges were involved, each responsible for grading a varying number of problems. Three of the most active judges had prior experience with evaluating LLM-generated proofs. One judge served as the coordinator, maintaining regular communication, tracking progress, and ensuring consistency and motivation across the group.
@@
Problems were drawn from top-tier national and international mathematics competitions, with the goal of capturing a balanced mix of correct and incorrect proofs. All problems were sourced from official materials. Non-English problems were translated using GPT-4.1 and manually verified for accuracy. When available, official ground-truth solutions were also extracted and provided to the judges as references.
@@
Proofs were generated using a set of state-of-the-art LLMs known for their strength in mathematical reasoning. Specifically, we used o4-mini and o3 from OpenAI (OpenAI, 2025), Gemini-2.5-Pro from Google (Google DeepMind, 2025), Grok-3-Mini from xAI (xAI, 2025), Qwen3-235B-A22B from Qwen (Qwen Team, 2025), and the latest version of DeepSeek-R1 from DeepSeek (Guo et al., 2025a). DeepSeek-R1 was released mid-way through dataset construction and replaced Grok-3-Mini thereafter. We designed the model prompt to clearly instruct models to generate full solutions, refining it through small-scale pilot tests. The final prompt is shown in Section˜E.1. All models were run with default parameters and a maximum generation length of 64,00064{,}000 tokens. If this limit was exceeded before a complete proof was generated, we re-sampled. For the MathArena subset (Balunović et al., 2025), we only retained solutions with a correct final answer, retrying generation if necessary. For PutnamBench (Tsoukalas et al., 2024), we appended the informal final answer (if present) to the problem statement to mirror the setup for formal models, allowing direct comparison between natural and formal proof outputs.
@@ 3.2 Grading Procedure
We built a custom web interface to facilitate efficient grading. A sample instance is available at https://judgeproofs.xyz/sample, with screenshots included in App.˜D. The interface displays the problem, reference solution (if available), anonymized model-generated proof, and grading form. Judges could mark the problem or solution as malformed (to filter such cases from the dataset), grade the proof, and annotate sentences from the model-generated proof with comments. Continuous feedback from judges helped us refine the interface over time.
@@
Judges were asked to label proofs as either correct or incorrect and provide written justification. Precise grading guidelines were critical to avoid inconsistencies in edge cases, such as minor omissions or shortcuts. To prevent overly strict grading, we clearly defined what level of omissions and frequency of mistakes were acceptable in a correct proof. The full guidelines are available at https://judgeproofs.xyz/instructions/, with a summary in App.˜D. These instructions were developed collaboratively with the judges and finalized after a pilot phase.
@@
After several hundred graded proofs, we introduced a new feature to support grading: an automatically generated summary of the proof using o4-mini. These summaries flagged potential issues—such as logical gaps or missing steps—based on a specially designed prompt (see Section˜E.2). Importantly, the model was instructed not to give a final verdict, only to indicate possible issues for human review. Judges reported that this significantly improved their efficiency and accuracy in detecting errors. To ensure the inclusion of these summaries did not bias our judges, we evaluated the agreement rate between o4-mini as a judge and human graders before and after their introduction. There was no significant difference in agreement, suggesting no introduction of bias into the grading process. However, in experiments involving best-of-n selection—where the LLM judge acts as a selection mechanism—we omitted these summaries to avoid any form of compounding bias in the evaluation.
@@
Problems were assigned to judges based on their mathematical background and availability. Judges not qualified to evaluate undergraduate-level problems were excluded from grading those problems.

### v2

@@ Abstract
In recent months, large language models (LLMs) have made significant progress in mathematical proof generation, but further advancement is hindered by the lack of a large-scale, high-quality dataset of human-evaluated proofs. While expensive to create, such a dataset is essential for driving improvements in training and addressing key open questions in the field of automated proof generation. Specifically, it remains unknown (1) how large the gap is between natural language and formal proof generation, (2) how final-answer accuracy relates to full proof correctness, and (3) how best-of-n selection strategies can affect proof quality. In this work, we present the Open Proof Corpus (OPC), a dataset comprising over 5,000 human-evaluated proofs produced by state-of-the-art LLMs. The OPC was specifically designed for broad applicability and downstream usage in proof generation research and is the first large dataset of LLM-generated solutions to problems from prestigious mathematics competitions such as the USAMO and IMO. Using the OPC, we address the open questions outlined above and provide new insights into LLMs’ strengths and limitations in mathematical reasoning. Finally, to showcase the utility of the OPC, we finetune an 8B-parameter model on the dataset, obtaining a model that matches Gemini-2.5-Pro, and performs close to the best model, GPT-5, on evaluating proof correctness.
@@ 1 Introduction
Large language models (LLMs) have recently achieved impressive progress in mathematical reasoning, obtaining top-competitor performance on various final-answer benchmarks such as AIME and HMMT (Balunović et al., 2025). However, growing evidence suggests that these benchmarks fail to capture the full range of mathematical capabilities, as they do not require models to produce proofs or detailed intermediate steps (Mahdavi et al., 2025; Guo et al., 2025b). Such step-by-step reasoning is critical for applications in theorem proving, mathematical research, and education.
@@
To address this, several evaluation efforts have taken place, revealing that LLMs significantly underperform on proof generation compared to existing final-answer benchmarks (Petrov et al., 2025; Mahdavi et al., 2025). Despite their value, these benchmarks are severely limited in their use for broader analysis and training. Specifically, they are small (Petrov et al., 2025; Guo et al., 2025b), use outdated models (Frieder et al., 2023), contain few correct proofs (Mahdavi et al., 2025), and are not open-sourced (Mahdavi et al., 2025; Guo et al., 2025b).
@@
Furthermore, key questions about proof generation capabilities remain unanswered. First, while it is widely claimed that there is a gap between final-answer performance and proof generation capabilities, this claim has not yet been supported by evaluating LLM-generated proofs on existing final-answer benchmarks. Second, despite recent advances in formal proof generation with Lean (Chen et al., 2025; Lin et al., 2025b), the performance gap between natural language and formal proof generation remains unclear. Third, the potential of best-of-n selection strategies to improve proof quality has not been explored.
Sample Problem: EGMO 2012, Problem 3 Level: High School Split: Generic Question: Solve over $\mathbb{R}$ the functional equation $f(yf(x+y)+f(x))=4x+2yf(x+y)$ Refer to caption Proof: To solve the functional equation …we conclude that this is the **only** solution. Refer to caption Grade (Incorrect): The model substitutes $x$ to be 1. All further ideas are based on the wrong assumption. Refer to caption
@@
Figure 1: Overview of the OPC and its conclusions. On the left, we show a typical sample, including a question from a high-quality mathematical competition, an LLM-generated proof, and a human judgment of proof correctness. On the right, we summarize the main conclusions from the OPC.
@@
To address these challenges, we introduce the Open Proof Corpus (OPC): a large-scale, human-validated dataset comprising over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems. As shown in Fig.˜1(a), each OPC sample includes (1) a problem from a high-quality mathematical competition such as the International Mathematical Olympiad, (2) a proof generated by a state-of-the-art LLM, and (3) a binary human evaluation of the proof’s correctness with feedback. The OPC was specifically designed for downstream usage in proof generation research, enabling both training and evaluation of LLMs on proof generation tasks. To address the open questions outlined earlier, the OPC includes problems from specific sources: problems from the PutnamBench (Tsoukalas et al., 2024) enable comparisons between formal and informal reasoning, while problems from MathArena (Balunović et al., 2025) support evaluation of proof correctness for problems with final answers.
@@
Despite the difficulty of problems in the OPC, state-of-the-art LLMs demonstrate surprisingly strong performance. For instance, o4-mini correctly solves almost $20\%$ of the problems in the IMO Shortlist, and the OPC generally consists of $43\%$ correct proofs. Furthermore, as shown in the top left of Fig.˜1(b), LLMs exhibit strong capabilities in evaluating proofs: GPT-5 achieves $90.8\%$ accuracy in judging proof correctness, which is on-par with human performance on this task. To showcase the utility of the OPC, we fine-tune R1-Qwen3-8B (Guo et al., 2025a) using GRPO (Shao et al., 2024) on the OPC, resulting in an open model that achieves $88.1\%$ judgment accuracy, close to top models like GPT-5 and outperforming the majority of frontier models.
@@
The OPC allows us to empirically resolve the open questions outlined above, with all conclusions shown in Fig.˜1(b). First, as shown on the top right, we find that natural language proof generation significantly outperforms formal proof generation: on PutnamBench, Gemini-2.5-Pro solves 4 times more problems than the best formal model, Goedel-Prover-V2 (Lin et al., 2025b). Second, as shown on the bottom left, we observe a substantial gap between final-answer accuracy and proof correctness. While Gemini-2.5-Pro loses only $8\%$ of its final-answer accuracy when proof correctness is required, o3 suffers a drop of almost $30\%$. Third, as shown on the bottom right, we find that best-of-n strategies yield large gains in performance. Interestingly, while standard best-of-n selection methods moderately improve accuracy from $26\%$ to $36\%$, a ranking approach inspired by (Liu et al., 2025) achieves the highest performance of $47\%$.
@@
  • A rigorous pipeline for generating and evaluating natural language proofs (Section˜3).
  • The OPC, a human-validated dataset comprising over 5,000 LLM-generated proofs (Section˜4).
  • Resolution of open questions in natural language proof generation using the OPC (Section˜5).
  • An open-source, 8B-parameter model fine-tuned on the OPC that achieves $88.1\%$ judgment accuracy, close to the best model on this task (Section˜5).
@@ 2 Related Work
We briefly discuss the relevant literature on mathematical reasoning benchmarks and datasets.
@@
Final-answer benchmarks evaluate models by comparing a final answer from the model’s output with a ground-truth answer, typically using rule-based parsers. With the rise of reasoning LLMs (Jaech et al., 2024), older benchmarks have become saturated (Cobbe et al., 2021; Lightman et al., 2024), and more recent ones are nearing saturation (Gao et al., 2025; He et al., 2024; Balunović et al., 2025). Only private benchmarks such as FrontierMath (Glazer et al., 2024) remain sufficiently challenging for the latest models. However, this benchmark does not require the generation of full proofs or detailed reasoning steps. Moreover, its private nature hinders reproducibility and broader community engagement, both key factors in driving progress.
@@
Another growing line of work involves training LLMs to generate formal proofs in languages such as Lean (de Moura and Ullrich, 2021) or Isabelle (Nipkow et al., 2002), which can then be automatically verified by these systems (Zheng et al., 2022; Tsoukalas et al., 2024; Yu et al., 2025). While this paradigm enables rigorous correctness checking, it typically requires models to be specifically finetuned for formal proof generation (Ren et al., 2025; Wang et al., 2025; Lin et al., 2025a). In contrast, state-of-the-art general-purpose models like GPT-5 and Gemini-2.5-Pro struggle with formal proof generation. As a result, formal reasoning currently does not make full use of the natural language capabilities of these general-purpose models, and, as we show in Section˜5, there remains a significant performance gap between formal and natural language proof generation, with the latter being substantially more effective.
@@
Two recent works have expanded this line of research. Sheng et al. (2025) focus specifically on inequality proofs and primarily emphasize the development of an LLM as a judge framework to mitigate the high cost of human evaluation. Guo et al. (2025b) highlight a notable gap between final-answer accuracy and the ability to generate correct proofs, a finding we confirm in Section˜5. However, their analysis stops at this observation and does not evaluate performance on an established final-answer benchmark. Further, both studies are limited in scale compared to the OPC and have not open-sourced their human-annotated proofs.
Automated proof grading with LLMs as judges
The growing capabilities of LLMs have led to the “LLM-as-a-judge” paradigm, which enables cost-effective evaluation of complex model outputs (Zheng et al., 2023; Gu et al., 2024; Chevalier et al., 2024). Recent work has extended this approach to mathematical proof evaluation (Sheng et al., 2025; Zhao et al., 2024; 2025). These studies report promising results, showing that LLMs can grade specific types of problems with reasonable reliability (Sheng et al., 2025; Zhao et al., 2024) and can provide useful feedback in educational settings (Zhao et al., 2024; 2025). This work builds on these insights, building a dataset of human-evaluated proofs to further improve LLM judging capabilities, as detailed in Section˜5.
@@
Several large-scale datasets have been developed to train LLMs on mathematical reasoning. One of the earliest, Li et al. (2024), compiled a large dataset of internet-sourced problems, including both final-answer questions and natural language proofs. However, it lacks LLM-generated proofs, human judgments, and examples of incorrect proofs. Other datasets have focused exclusively on final-answer tasks (Albalak et al., 2025; He et al., 2025; Moshkov et al., 2025), offering limited support for training or evaluating proof generation. Finally, Zhang et al. (2025) introduced a dataset of both valid and invalid problem statements, each accompanied by LLM-generated proofs. While this ensures that incorrect proofs exist for invalid statements, the dataset does not include human evaluation of proofs or other information on proof validity.
@@ 3 Methodology
Accurately evaluating LLM-generated proofs is a challenging task. Models often make difficult-to-detect errors, and they rarely acknowledge when they cannot solve a problem (see Section˜5). In this section, we outline the methodology used to create the OPC, with particular emphasis on the complexities of evaluating LLM-generated proofs and our efforts to maximize dataset size. Since human judges cannot reasonably spend hours studying each problem, we developed a pipeline to support efficient grading. This pipeline consisted of three main components: problem and judge preparation (Section˜3.1), the grading procedure (Section˜3.2), and monitoring and validation (Section˜3.3). The dataset was constructed over a period of four weeks, involving $13$ expert judges and generating over 5,0005{,}000 human-validated proofs.
@@ 3.1 Problem and Judge Preparation
Judges were selected from among former IMO participants or individuals who reached the final stages of IMO selection in their countries. We contacted each judge personally to ensure they were qualified and motivated. A total of $13$ judges were involved, each responsible for grading a varying number of problems. Three of the most active judges had prior experience with evaluating LLM-generated proofs. One judge served as the coordinator, maintaining regular communication, tracking progress, and ensuring consistency and motivation across the group.
@@
Problems were drawn from top-tier national and international mathematics competitions, such as the USAMO and IMO, to capture a balanced mix of correct and incorrect proofs. In particular, we included competitions based on two informal criteria: (1) the competition is well-known and produces high-quality problems, and (2) the difficulty level of the competition aligns with our target of roughly 50% model accuracy. All problems were sourced from official materials. Non-English problems were translated using GPT-4.1 and manually verified for accuracy by the coordinator. When available, official ground-truth solutions were also extracted and provided to the judges as references.
Throughout the annotation process, model performance was actively monitored to ensure that the selected problems remained appropriately challenging. For instance, more problems from international competitions were added when initial results indicated that models were performing very well ($\approx 65\%$) on national-level problems. Each day, problem prioritization was adjusted based on ongoing performance metrics, judge availability, and progress towards the specific conclusions we aimed to draw from the dataset.
@@
Proofs were generated using a set of state-of-the-art LLMs known for their strength in mathematical reasoning. Specifically, we used o4-mini and o3 from OpenAI (OpenAI, 2025), Gemini-2.5-Pro from Google (Team, 2025), Grok-3-Mini from xAI (xAI, 2025), Qwen3-235B-A22B from Qwen (Qwen Team, 2025), and the latest version of R1 from DeepSeek (Guo et al., 2025a). R1 was released mid-way through dataset construction and replaced Grok-3-Mini thereafter. We designed the model prompt to clearly instruct models to generate full solutions, refining it through small-scale pilot tests. The final prompt is shown in Section˜I.1. All models were run with default parameters and a maximum generation length of 64,00064{,}000 tokens. For the MathArena subset (Balunović et al., 2025), we only retained solutions with a correct final answer, retrying generation if necessary. For PutnamBench (Tsoukalas et al., 2024), we appended the informal final answer (if present) to the problem statement to mirror the setup for formal models, allowing direct comparison between natural and formal proof outputs.
@@ 3.2 Grading Procedure
We built a custom web interface to facilitate efficient grading. A sample instance is available in our supplementary material, with screenshots included in App.˜H. The interface displays the problem, reference solution (if available), anonymized model-generated proof, and grading form. Judges could mark the problem or solution as malformed (to filter such cases from the dataset), grade the proof, and annotate sentences from the model-generated proof with comments. Continuous feedback from judges helped us refine the interface over time.
@@
Judges were asked to label proofs as either correct or incorrect and provide written justification. Precise grading guidelines were critical to avoid inconsistencies in edge cases, such as minor omissions or shortcuts. To prevent overly strict grading, we clearly defined what level of omissions and frequency of mistakes were acceptable in a correct proof. The full guidelines are available in our supplementary material, with a summary in App.˜H. These instructions were developed collaboratively with the judges and finalized after a pilot phase (see Section˜3.3).
@@
After several hundred graded proofs, we introduced a new feature to support grading: an automatically generated summary of the proof using o4-mini. These summaries flagged potential issues, such as logical gaps or missing steps, based on a specially designed prompt (see Section˜I.2). Importantly, the model was instructed not to give a final verdict. Judges reported that this significantly improved their efficiency and accuracy in detecting errors. To ensure the inclusion of these summaries did not bias our judges, we evaluated the agreement rate between o4-mini as a judge and human graders before and after their introduction. There was no significant difference in agreement, indicating that no bias was introduced into the grading process. However, in experiments involving best-of-n selection, where the LLM judge acts as a selection mechanism, we omitted these summaries to avoid any form of compounding bias in the evaluation.
@@
Problems were assigned to judges based on background and availability. Judges not qualified to evaluate undergraduate-level problems were excluded from grading them. To ensure a balanced workload, we monitored grading progress and dynamically reassigned problems as needed.